a 0.0585 kg tennis ball is twirled at the end of a thread in a horizontal circle with radius of 0.75m with a height of 2m above the ground. the thread breaks and the balls travels a horizontal distance of 18.4 m before bouncing off the ground. what is the tension on the thread before it snapped?
1
Expert's answer
2019-11-04T10:39:48-0500

For calculating the velocity just before breaking of string


time of flight = "\sqrt{\frac{2h}{g}}=\sqrt\frac{4}{10}=0.635\ sec"


horizontal range = "vt=0.635v=18.4"

"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ v=28.98\ m\ sec^{-1}"

now, tension,T

"T =\frac{mv^2}{r}=\frac{0.0585\times(28.98)^2}{0.75}=65.5\ N"